OpenStudy (anonymous):

A 2m tall woman is running at a rate of 2 m/s when she spots a bird in a tree 6m off the ground. How fast is the angle of elevation to the bird increasing when she is 3m from the base of the tree?

OpenStudy (amistre64):

can you define an equation to model the situation?

OpenStudy (amistre64):



OpenStudy (amistre64):

\[tan(\alpha)=\frac4d\]

OpenStudy (amistre64):

take the derivaitve \[tan(\alpha)=\frac4d\] \[sec^2(\alpha)~\alpha'=-\frac{4d'}{d^2}\]

OpenStudy (amistre64):

we know d, we can readily calculate a, and we know d'

OpenStudy (amistre64):

\[\alpha'=-\frac{4(-2)}{3^2}~(\frac{3}{5})^2\] -2 since they are running towards the tree, or play with the sign

OpenStudy (amistre64):

the closer you get, the faster yo have to look up a a fixed point i beleive